The cross section of the iron conductor is a square of 1.2 mm. What is the length of the conductor if its resistance is 10 Ω? ρ = 10-7 Ωm.




1
Expert's answer
2021-05-31T08:39:54-0400

The resistance is given as follows:


"R = \rho\dfrac{l}{S}"

where "\rho = 10^{-7}\Omega\cdot m" is the resestivity of the material, "l" is the length of the conductor, and "S = 1.2^2mm^2 = 1.44mm^2 = 1.44\times 10^{-6}m^2" is the area of its cross section. Expressing "l" and substituting "R= 10\Omega", obtain:


"l = \dfrac{RS}{\rho}\\
l = \dfrac{10\cdot 1.44\times 10^{-6}}{10^{-7}} = 144m"

Answer. 144 m.